# Classification Diffusion Models: Revitalizing Density Ratio Estimation

### v2

@@ C.1 Architectures
For fair comparison, for each dataset we used the same architecture for our method and for DDMs. As baslines we took the pre-trained model for CelebA $64\times 64$ from the DDIM Official Github [38] and the EMA pre-trained model for CIFAR-10 from the pytorch diffusion repository, who converted the pre-trained model from the Official DDPM implementation from tensorflow to pytorch.
For conditional CIFAR-10 we trained the DDM model by ourselves because there exist no pre-trained models for CIFAR-10 capable of handling CFG. We used the same architecture from [19] for both models and trained them for the same number of iterations (more details are in App. C.2.2). To condition the model on the class label, we learned an embedding for each class using nn.Embedding and injected it at all points where the time embedding was injected in the original architecture (for DDM we added it to the time embedding and for CDM we replaced the time embedding by the class embedding).
@@ C.2.2 Unconditional and Conditional CIFAR-10
We trained the model for 500k iterations with a learning-rate of $2\cdot 10^{-4}$. We started with a linear warmup of 5k iterations and reduced the learning rate by a factor of 10 after every 200k iterations. The typical value of the CE loss after convergence was $\sim 4.4$ while the MSE loss was $\sim 0.03$ so we chose to give the CE loss a weight of $0.001$ to maintain values at the same order of magnitude. In addition, we used EMA with a factor of 0.9999, as done in the baseline model.
@@ C.6 The Addition of The Timesteps 00 and $T+1$
As outlined in Sec. 3, we introduce an additional timesteps ${\bm{x}}_{0}$, ${\bm{x}}_{T+1}$ that are not present in DDMs. This inclusion is fundamental to our method formulation. Importantly, this addition does not alter the number of sampling steps, as we continue initiating the reverse process from ${\textnormal{t}}=T$ and finish it with the denoised result of ${\textnormal{t}}=1$, following the approach in DDM. During training, we employ both MSE and CE losses for all timesteps, while the first timestep ${\textnormal{t}}=0$ and the last timestep ${\textnormal{t}}=T+1$ are trained without MSE loss.
Appendix D Toy Example Log Likelihood
Figure 7: Analytical Example of Log Likelihood. The illustrated densities represent uniform distributions $U\left[-\frac{\gamma}{2},\frac{\gamma}{2}\right]^{3\times 32\times 32}$ for each $\gamma$ value in the set $\{0.25,0.5,1,2,3\}$. Samples from these distributions correspond to images of iid uniform noise. The red points on the graph indicate the estimated likelihood results obtained from CDM models trained on samples from each distribution. The alignment of these points along the diagonal serves as evidence supporting our likelihood estimation as per Theorem 3.2.
In Fig. 7 we validated our efficient likelihood computation by applying it to toy examples with known densities, allowing us to compute the analytical likelihood and verify that our computations align with theoretical expectations. We chose the uniform distributions $p_{\gamma}=U\left[-\frac{\gamma}{2},\frac{\gamma}{2}\right]^{3\times 32\times 32}$ for each $\gamma$ value in the set $\{0.25,0.5,1,2,3\}$ and trained CDM individually for each density following the procedure outlined in Algorithm 1. These distributions correspond to $32\times 32$ color images of iid uniform noise. The probability density function (pdf) of the uniform distribution $U\left[-\frac{\gamma}{2},\frac{\gamma}{2}\right]$ is given by
@@ Appendix D Toy Example Log Likelihood
The logarithm of the pdf $p_{\gamma}({\bm{x}})$ is given by $-d\ln(\gamma)$. The log-likelihood is the expectation of $\ln p_{\gamma}({\bm{x}})$. Since $\ln p_{\gamma}({\bm{x}})$ is constant, its expectation is that constant. Therefore, the log-likelihood normalized by $d$ is $-\ln(\gamma)$.
@@ E.1 Analysis of the Effect of Training with Different Losses
(b) Confusion matrix CDM(unif.) model evaluated on unconditional CIFAR-10 $\mathbf{32\times 32}$ using the scheduler from TDR. The colors indicate the probabilities. In contrast to DDPM noise scheduler, with TDR noise scheduler, the classification difficulty is preserved across all timesteps.

### v3

@@ C.1 Architectures
For a fair comparison, for each dataset we used the same architecture for our method and for DDMs. As baslines we took the pre-trained model for CelebA $64\times 64$ from the DDIM Official Github [39] and the EMA pre-trained model for CIFAR-10 from the pytorch diffusion repository, who converted the pre-trained model from the Official DDPM implementation from tensorflow to pytorch.
For conditional CIFAR-10 we trained the DDM model by ourselves because there exist no pre-trained models for CIFAR-10 capable of handling CFG. We used the same architecture from [19] for both models and trained them for the same number of iterations (more details are in App. C.2.2). To condition the model on class labels, we learned an embedding for each class using nn.Embedding and injected it at all points where the time embedding was originally applied. In the case of DDM, we added the class embedding to the time embedding, while for CDM, we replaced the time embedding with the class embedding.
@@ C.2.2 Unconditional and Conditional CIFAR-10
We trained the model for 500k iterations with a learning rate of $2\cdot 10^{-4}$. We started with a linear warmup of 5k iterations and reduced the learning rate by a factor of 10 after every 200k iterations. The typical value of the CE loss after convergence was $\sim 4.4$ while the MSE loss was $\sim 0.03$ so we chose to give the CE loss a weight of $0.001$ to maintain values at the same order of magnitude. In addition, we used EMA with a factor of 0.9999, as done in the baseline model.
@@
The CDM(OT) model was trained with the same hyperparameters, except for the learning rate, which was set to $1\cdot 10^{-4}$. Additionally, we trained the model without a learning rate schedule for 1M iterations, as the NLL continued to decrease beyond 500k iterations.
@@ C.6 The Addition of The Timesteps 00 and $T+1$
As outlined in Sec. 3, we introduce additional timesteps, corresponding to ${\bm{x}}_{0}$, ${\bm{x}}_{T+1}$, which are not present in DDMs. This inclusion is fundamental to our method formulation. Importantly, this addition does not alter the number of sampling steps, as we continue initiating the reverse process from ${\textnormal{t}}=T$ and finish it with the denoised result of ${\textnormal{t}}=1$, following the approach in DDM.
Appendix D Validating the Log Likelihood Computation on Toy Examples
We validate our efficient likelihood computation by applying it to toy examples with known densities, allowing us to compute the analytical likelihood and verify that our computations align with theoretical expectations.
D.1 Images With Independent Pixels
Figure 7: Log likelihood computation using CDM for toy problems possessing closed form expressions. The left subplot corresponds to the densities from Sec. D.1, while the right subplot corresponds to the densities from Sec. D.2. The red points on the graphs correspond to differnet $\gamma$ values. The alignment of these points along the diagonal provides evidence supporting our likelihood estimation as stated in Theorem 3.2.
We start by experimenting with the uniform distributions $p_{\gamma}=U\left[-\frac{\gamma}{2},\frac{\gamma}{2}\right]^{3\times 32\times 32}$ with $\gamma\in\{0.25,0.5,1,2,3\}$. We trained a CDM separately for each of these densities following the procedure outlined in Algorithm 1. These distributions correspond to $32\times 32$ color images of iid uniform noise. The probability density function (pdf) of the uniform distribution $U\left[-\frac{\gamma}{2},\frac{\gamma}{2}\right]$ is given by
@@ D.1 Images With Independent Pixels
The logarithm of the pdf $p_{\gamma}({\bm{x}})$ is given by $-d\ln(\gamma)$. The log-likelihood is the expectation of $\ln p_{\gamma}({\bm{x}})$. Since $\ln p_{\gamma}({\bm{x}})$ is constant, its expectation is that constant. Therefore, the log-likelihood normalized by $d$ is $-\ln(\gamma)$. In the left subplot of Fig. 7 we compare the analytical log likelihood computed using (32) with the estimated log likelihood by our model.
D.2 Images With Correlated Pixels
To further validate our likelihood calculation, we tested it on toy examples of images with correlated pixels. Specifically, we used images of size $3\times 32\times 32$ sampled from multivariate Gaussian distributions with $\mu=0$ and $\Sigma\neq I$. The normalized log likelihood per dimension for a multivariate Gaussian is given by
$-0.5\cdot\left(\ln(2\pi)+1\right)-\frac{\ln(|\Sigma|)}{2d}.$ (32)
We defined a sequence of distributions with $\Sigma=\sqrt{1-\gamma}\,\Sigma_{\text{CIFAR-10}}+\sqrt{\gamma}\,I$, where $\Sigma_{\text{CIFAR-10}}$ is the empirical covariance matrix of the CIFAR-10 dataset, and $\gamma\in\{0,10^{-5},10^{-3},10^{-1}\}$. In the right subplot of Fig. 7, we compare the analytical log likelihood computed using (32) with the estimated log likelihood by our model.
@@ E.1 Analysis of the Effect of Training with Different Losses
(b) Confusion matrix CDM(unif.) model evaluated on unconditional CIFAR-10 $\mathbf{32\times 32}$ using the scheduler from TRE. The colors indicate the probabilities. In contrast to DDPM noise scheduler, with TDR noise scheduler, the classification difficulty is preserved across all timesteps.